# Finite temperature dopant-induced spin reorganization explored via tensor networks in the two-dimensional $t$-$J$ model

### v2

@@ Abstract
Doped Mott insulators host intertwined spin-charge phenomena that evolve with temperature and can culminate in stripe order or superconductivity at low temperatures. The two-dimensional $t$-$J$ model captures this interplay yet finite-temperature, infinite-size calculations remain difficult. Using purification represented by a tensor network – an infinite projected entangled-pair state (iPEPS) ansatz – we simulate the $t$-$J$ model at finite temperature directly in the thermodynamic limit, reaching temperatures down to one tenth of the hopping rate and hole concentrations up to one quarter of the lattice sites. Beyond specific heat, uniform susceptibility, and compressibility, we introduce dopant-conditioned multi-point correlators that map how holes reshape local exchange. Nearest-neighbor hole pairs produce a strong cooperative response that reinforces antiferromagnetism on the adjacent parallel bonds, and single holes weaken nearby antiferromagnetic bonds; $d$-wave pairing correlations remain short-ranged over the same window. These results provide experiment-compatible thermodynamic-limit benchmarks and establish dopant-conditioned correlators as incisive probes of short-range spin-texture reorganization at finite temperature.
INTRODUCTION
Strongly correlated electron systems on a two-dimensional (2D) lattice exhibit a spectacular variety of collective behaviors, all rooted in three simple ingredients: the tendency of electrons to hop between sites (kinetic energy), their magnetic superexchange interactions, and thermal fluctuations. In materials ranging from the cuprates and nickelates to layered ruthenates and twisted-bilayer graphene, doping a Mott-insulating antiferromagnet gives rise to intertwined spin and charge textures often organizing into stripes [1, 2] and to unconventional superconductivity with $d$-wave or other exotic pairing symmetries [3]. As temperature rises, these same systems enter the enigmatic pseudogap regime, characterized by a partial suppression of the electronic density of states [4], and on further doping the strange-metal regime, where transport violates conventional Fermi-liquid rules [5].
@@ INTRODUCTION
Inspired by these advances, we use purification [39, 40, 41] with infinite projected entangled pair states (iPEPS) [42, 43, 44, 45] ansatz to study the two dimensional square lattice $t$-$J$ model at finite temperature directly in the thermodynamic limit, with significant technical improvements over the technique applied to a previous study on the Fermi-Hubbard model [46]. By reaching temperatures as low as $1/10$ of tunneling energy and doping up to $25\%$, we first map out bulk thermodynamics — specific heat, uniform spin susceptibility, and charge compressibility — to pinpoint the onset of strong spin-charge interplay. Guided by these macroscopic signatures, we then compute conditional three-point and higher order spin-charge and fermionic correlators to reveal how holes reorganize their magnetic environment. Our results provide quantitative benchmarks for both quantum-gas experiments and other numerical approaches trying to reach the thermodynamic limit, and shed new light on the microscopic mechanisms driving emergent phenomena in doped Mott insulators.
Results
@@
where the operator $\tilde{c}_{i\sigma}$ projects out double occupancy, $\mu$ is the chemical potential, and $n_{i}$ is the electron number operator. In all numerical simulations, we control doping by tuning $\mu$ and measure the filling $n$. Unless stated otherwise, all quantities are plotted vs. temperature $T$ in units of the hopping $t=1$. The filling $n$ runs from $1$ (half-filled) down to $\approx 0.75$. In the following we set $J=0.5$. Please refer to Appendix A to see how doping depends on chemical potential and temperature $T$. Fig. 1 illustrates the dynamics of the $t$-$J$ model and the principal physics in the article. Panel A is a classical sketch of the doped $t$-$J$ model lattice: spins (up: red; down: blue) occupy some lattice sites, other lattice sites have holes denoted by empty gray circles. Here, allowed moves are spin hops into a neighboring hole occupied site, and any move that would create double occupancy is forbidden (marked with a red $\times$); bonds connecting two occupied sites with opposite spins are highlighted with a green halo to indicate that antiparallel occupied bonds have superexchange contribution $\sim J$, while bonds touching a hole or like spins carry no exchange contribution. Panel B shows how dopants reorganize the local magnetism of an antiferromagnetic background. The three scenes correspond to the unconditioned AFM background (center), one fixed hole (left), and two neighboring fixed holes (right). A single hole (a magnetic polaron) disturbs nearby AFM links and produces a short-range ferromagnetic skew around it, whereas two adjacent holes moving together break fewer AFM bonds than two isolated holes and therefore cooperatively stabilize an antiferromagnetic segment along the plaquette edge parallel to the pair.
@@
The method of choice here is purification within an infinite projected entangled-pair state ansatz, which has been previously successfully applied to a variety of problems [41, 48, 49, 50, 51, 52, 53]. The purification is evolved in imaginary time from $T=\infty$ to a given temperature. The ansatz uses a two-site checkerboard with $U(1)\times U(1)$ symmetry. We optimize imaginary-time gates using a neighborhood tensor update [45, 54, 55] with an environment assisted initial truncation [46]. We evaluate observables with a multiplet-aware extension of the corner transfer matrix renormalization group [56, 57, 10]. Working directly in the thermodynamic limit avoids finite-size bias. The correlations are then computed using the zipper algorithm [58]. Please check the Methods section for details. All simulations were performed with the YASTN library [59, 60].
Thermodynamics
@@
Fig. 2 summarizes bulk thermodynamics of the square lattice $t$-$J$ model in the thermodynamic limit. We focus especially on the low-to-intermediate temperature range $0.1\lesssim T\lesssim 0.5$, where short-range AFM order and emergent pseudogap features become prominent [61]. Fig. 2(A) shows the specific heat $C_{v}(T)=\partial_{T}{\langle H\rangle}$, which exhibits the expected spin-entropy peak associated with the formation of short-range antiferromagnetism. At (or very near) half filling ($n=1$), a broad maximum appears at $T\sim J$, reflecting the release of entropy when nearest-neighbor singlets proliferate and spin correlations grow. Upon increasing doping 1n1-n, the peak diminishes in height and shifts to higher $T$, consistent with mobile holes breaking AFM bonds, reducing the spin stiffness, and shortening the antiferromagnetic length. In the high temperature window, $T\geq 1$, the inset reveals the anticipated linear in 1n1-n behavior of $C_{v}$ from high $T$ expansion: energy fluctuations are dominated by uncorrelated spins and holes.
@@
In this work, we set $h=0.1$ and $h=0.01$. The results of these fields match with each other quantitatively, proving that we are in the weak-field regime. For weak-doping regime (e.g., $n\sim 0.95$), $\chi(T)$ develops a clear maximum at a temperature $T^{*}$ (see the inset of Fig. 2(B)), which tracks the build-up of strong but short-ranged AFM correlations: cooling below $T^{*}$ suppresses spin fluctuations as local singlets form; heating above $T^{*}$ melts the AFM background into a Curie-like regime. As doping increases, the maximum weakens and shifts to lower $T$. We thus identify $T^{*}$ as a pseudogap crossover scale set by short-range [62, 27]. Its evolution mirrors the suppression of the $C_{v}$ peak and anchors the microscopic reorganization discussed later. In Appendix B, we compare our iPEPS results for specific heat and magnetic susceptibility against METTS using matrix product states applied to cylinders and have appreciable agreement.
@@
Fig. 2(C) shows the isothermal compressibility $\kappa(\mu,T)=(\partial n/\partial\mu)_{T}$ together with filling $n(\mu)$ (inset). At high temperature, the response is nearly featureless, as expected for a classical mixture of spins and holes with weak correlations. Upon cooling to about $T\sim 0.3$ range, $\kappa$ develops a small peak as a function of $\mu$ (or equivalently $n$). This emerges in the same doping window where $\chi(T)$ exhibits a pronounced maximum. Two competing trends shape $\kappa$ here. Close to half filling, Mott constraints suppress charge fluctuations and reduce $\kappa$; moving away from half filling, the growth of short-range AFM domains lowers the energy cost of rearranging charge locally (e.g., by accommodating holes so as not to destroy the background AFM), mildly enhancing $\kappa$. The net result is a small but robust peak which grows as a function of decreasing temperature, consistent with a crossover — not a phase transition — in two dimensions at finite $T$. This intermediate-$T$ peak of $\kappa$ versus doping mirrors the cluster-DMFT Widom-line phenomenology of the doped Mott insulator, where $\kappa$ and related thermodynamic derivatives reach maxima above a pseudogap–metal critical endpoint and thus track the pseudogap boundary [63, 64, 65]. This phenomenon has also been tied to charge clustering [28] by one of the authors, and our METTS snapshots on a cylindrical lattice in Appendix D show the same growth of fluctuating hole clusters without macroscopic phase separation. There, we detail the analysis: we flag hole-rich sites via an adaptive threshold, group them into nearest-neighbor connected components on a cylinder of length $24$ and width $4$ with periodic boundary conditions, and compile density-weighted cluster-size histograms resolved by total hole mass. As the temperature is lowered, the probability for having clusters of a certain size shifts toward larger but still finite clusters. Our compressibility values also agree qualitatively with the finite-temperature Lanczos benchmarks of Ref. [23]. A recent Fermi–Hubbard quantum simulator study similarly mapped a pseudogap phase diagram via a doping-tuned compressibility maximum and additionally momentum-selective depletion of low-energy spectral weight, revealing a line of thermodynamic anomalies and suggesting a link to charge order [38].
Higher Order Correlators
@@
In Fig. 3, we evaluate the spin-spin-hole 3P correlators for bonds at distance $d=1$ or $\sqrt{2}$. In a Néel background the unconditional correlator $\langle B_{ab}\rangle$ is negative for nearest neighbours (d=1d{=}1) and positive for diagonals ($d{=}\sqrt{2}$). Hence a red (positive) $C^{(3)}$ on d=1d{=}1 means the AFM bond is weakened (made less negative) by the hole, whereas a blue (negative) $C^{(3)}$ on $d{=}\sqrt{2}$ means the originally positive diagonal correlation is suppressed, both indicating a local FM effect of the hole. At high to intermediate temperatures ($T\geq 1/4$), the correlation maps are short-ranged across $n=0.85$, $0.9$ and $0.95$, but grow upon cooling and with increasing filling, as shown in Fig. 3(A). At lower temperature $T=1/10$, the correlator extends over the whole $7\times 7$ window for all fillings with a much slower decay. Figs. 3(B), (C), and (D) show these general tendencies in more detail.
@@
In Fig. 4 we contrast the single-dopant L-geometry $C^{(3)}_{L}$ with two different two-dopant conditionings: the well separated $T$ configuration $C^{(4)}_{T}$, and the nearest-neighbor plaquette configuration $C^{(4)}_{\parallel}$ as a function of the doping 1n1-n. For dopings $0.00\lesssim 1-n\lesssim 0.18$, both $C^{(3)}_{L}$ and $C^{(4)}_{T}$ are positive, then cross zero near $n=0.8$ and become negative upon further doping. Moreover, the near-additivity $C^{(4)}_{T}\simeq 2C^{(3)}_{L}$ holds on both sides of the crossing within the plotted scatter, indicating weak interaction between the well-separated dopants. Interpreting the sign relative to the background bond value, the positive regime signals a local ferromagnetic skewing around a dopant (AFM bonds made less negative), whereas beyond filling $n\sim 0.8$, the softened AFM background (reduced spin stiffness) leads the conditioned response to reinforce the residual AFM correlations (AFM bonds made more negative). By contrast, $C^{(4)}_{\parallel}$ is negative across the entire range and grows in magnitude away from half filling (increased doping 1n1-n) and on cooling, consistent with the following mechanism: two adjacent holes eliminate seven, rather than eight, JJ–links, lowering the exchange cost and cooperatively enforcing AFM on the parallel edge of the plaquette. This response, at finite $T$, provides a microscopic route toward incipient charge segregation marked by enhanced compressibility (as shown in Fig. 2(C)); upon further cooling, it can act as a precursor to stripe formation where holes line up to minimize exchange frustration while retaining kinetic energy. Finally, the doping where $C^{(3)}_{L}$ and $C^{(4)}_{T}$ change sign coincides with the regime where the uniform susceptibility peak fades in thermodynamics (as shown in Fig. 2(B)), reinforcing the link between the weakening spin stiffness and the reorganization of local spin textures.
@@
In Appendix C, we show results for two other kinds of correlators: the five-point correlator consisting of four spin operators on the nearest-neighbor sites surrounding a central hole projector, and the electron momentum distribution function.
Discussion
We access the thermodynamic limit of the two-dimensional $t$-$J$ model down to $T\simeq 0.1$ and up to hole doping $1-n\lesssim 0.25$, using purification-iPEPS with a neighborhood tensor update and multiplet-aware CTMRG to stably evaluate bulk thermodynamics and hole-conditioned multi-point observables. The picture that emerges is that cooling builds short-ranged antiferromagnetism, single holes dress into magnetic polarons that locally skew nearest-neighbor exchange, and nearest-neighbor hole pairs cooperate to reduce exchange frustration on the parallel plaquette edge. Thermodynamically we see a broad spin-entropy feature in $C_{v}(T)$, a pseudogap-like maximum in the uniform susceptibility that shifts and fades with doping, and a shallow compressibility bump versus $\mu$ upon cooling. The $d$-wave pairing remains short-ranged over the same window, so magnetic reorganization precedes any superconducting coherence in our parameter set. These trends dovetail with the latest quantum-gas microscope evidence for a pseudogapped metal in the Hubbard model, where saturation of magnetic susceptibility versus temperature [36] and compressibility maximum versus doping develops upon cooling enabling a pseudogap phase diagram from thermodynamics [38]. Our conditional correlators supply a concrete short-distance mechanism for the thermodynamic anomaly and suggest why enhanced $\kappa$ can arise without macroscopic phase separation at finite $T$, and they provide thermodynamic limit benchmarks for cold atoms experiments. Our two-site checkerboard ansatz cannot host stripes or any longer ranged spin-charge order; larger unit cells will be needed to capture symmetry breaking and access the stripe regime. Taken together, the thermodynamic benchmarks plus hole-conditioned correlators show that — at experimentally relevant temperature and dopings — the finite temperature reorganization of the ttJJ model is predominantly magnetic and local, while superconducting correlations remain short-ranged.
Methods

### v3

@@ Abstract
We study the two-dimensional ttJJ model at finite temperature directly in the thermodynamic limit using purification represented by an infinite projected entangled-pair state (iPEPS). We reach temperatures down to $T/t=0.1$ and hole concentrations up to $1-n\simeq 0.25$, and provide benchmark thermodynamic-limit results for the specific heat, uniform susceptibility, and charge compressibility. We identify a susceptibility maximum $T^{\ast}$ that tracks the buildup of short-range antiferromagnetism and a shallow compressibility enhancement upon cooling in the same doping window. To expose the underlying microscopic mechanism, we introduce dopant-conditioned multi-point correlators that quantify how holes reorganize nearby exchange: single holes weaken adjacent antiferromagnetic bonds, while nearest-neighbor hole pairs produce a cooperative response that reinforces antiferromagnetism on the parallel plaquette edge. Over the same parameter window, $d$-wave pairing correlations remain short-ranged. These results provide experiment-compatible thermodynamic-limit benchmarks and establish dopant-conditioned correlators as incisive probes of finite-temperature spin-texture reorganization in doped Mott insulators.
I INTRODUCTION
Strongly correlated electron systems on a two-dimensional (2D) lattice exhibit a spectacular variety of collective behaviors, all rooted in three simple ingredients: the tendency of electrons to hop between sites (kinetic energy), their magnetic superexchange interactions, and thermal fluctuations. In materials ranging from the cuprates and nickelates to layered ruthenates and twisted-bilayer graphene [36, 45, 49, 6, 5], doping a Mott-insulating antiferromagnet gives rise to intertwined spin and charge textures often organizing into stripes [25, 36] and to unconventional superconductivity with $d$-wave or other exotic pairing symmetries [44]. As temperature rises, these same systems enter the enigmatic pseudogap regime, characterized by a partial suppression of the electronic density of states [73], and on further doping the strange-metal regime, where transport violates conventional Fermi-liquid rules [55].
@@ I INTRODUCTION
Inspired by these advances, we use purification [74, 15, 16] with infinite projected entangled pair states (iPEPS) [35, 12, 17, 21] ansatz to study the two dimensional square lattice $t$-$J$ model at finite temperature directly in the thermodynamic limit, with significant technical improvements over the technique applied to a previous study on the Fermi-Hubbard model [66]. In this paper we make three contributions. First, we provide thermodynamic-limit finite-$T$ benchmarks for the square-lattice ttJJ model down to $T/t=0.1$ across dopings up to $1-n\simeq 0.25$. Second, we extract and track the susceptibility maximum $T^{\ast}$ and a shallow compressibility enhancement upon cooling, which together delineate the regime of strongest short-range spin–charge interplay in our data. Third, we introduce dopant-conditioned three- and four-point correlators that directly visualize how single holes and nearest-neighbor hole pairs reorganize local exchange, and we relate these microscopic responses to the thermodynamic anomalies.
II Model
@@
where the operator $\tilde{c}_{i\sigma}$ projects out double occupancy, $\mu$ is the chemical potential, and $n_{i}$ is the electron number operator. In all numerical simulations, we control doping by tuning $\mu$ and measure the filling $n$. Unless stated otherwise, all quantities are plotted vs. temperature $T$ in units of the hopping $t=1$. The filling $n$ runs from $1$ (half-filled) down to $\approx 0.75$. In the following we set $J=0.5$.
@@
Fig. 1 illustrates the dynamics of the $t$-$J$ model and the principal physics in the article. Panel A is a classical sketch of the doped $t$-$J$ model lattice: spins (up: red; down: blue) occupy some lattice sites, other lattice sites have holes denoted by empty gray circles. Here, allowed moves are spin hops into a neighboring hole occupied site, and any move that would create double occupancy is forbidden (marked with a red $\times$); bonds connecting two occupied sites with opposite spins are highlighted with a green halo to indicate that antiparallel occupied bonds have superexchange contribution $\sim J$, while bonds touching a hole or like spins carry no exchange contribution. Panel B shows how dopants reorganize the local magnetism of an antiferromagnetic background. The three scenes correspond to the unconditioned AFM background (center), one fixed hole (left), and two neighboring fixed holes (right). A single hole (a magnetic polaron) disturbs nearby AFM links and produces a short-range ferromagnetic skew around it, whereas two adjacent holes moving together break fewer AFM bonds than two isolated holes and therefore cooperatively stabilize an antiferromagnetic segment along the plaquette edge parallel to the pair.
@@
The method of choice here is purification within an infinite projected entangled-pair state ansatz, which has been previously successfully applied to a variety of problems [16, 42, 18, 19, 33, 56, 52]. The purification is evolved in imaginary time from $T=\infty$ to a given temperature. The ansatz uses a two-site checkerboard with $U(1)\times U(1)$ symmetry. We optimize imaginary-time gates using a neighborhood tensor update [21, 22, 48] with an environment assisted initial truncation [66]. We evaluate observables with a multiplet-aware extension of the corner transfer matrix renormalization group [53, 51, 13]. Working directly in the thermodynamic limit avoids finite-size bias. The correlations are then computed using the zipper algorithm [67]. Please check Appendix A for methodological and numerical details. All simulations were performed with the YASTN library [62, 61].
III Thermodynamics
@@
Fig. 2 summarizes bulk thermodynamics of the square lattice $t$-$J$ model in the thermodynamic limit. We focus especially on the low-to-intermediate temperature range $0.1\lesssim T\lesssim 0.5$, where short-range AFM order and emergent pseudogap features become prominent [58]. Fig. 2(A) shows the specific heat $C_{v}(T)=\partial_{T}{\langle H\rangle}$, which exhibits the expected spin-entropy peak associated with the formation of short-range antiferromagnetism. At (or very near) half filling ($n=1$), a broad maximum appears at $T\sim J$, reflecting the release of entropy when nearest-neighbor singlets proliferate and spin correlations grow. Upon increasing doping 1n1-n, the peak diminishes in height and shifts to higher $T$, consistent with mobile holes breaking AFM bonds, reducing the spin stiffness, and shortening the antiferromagnetic length. In the high temperature window, $T\geq 1$, the inset reveals the anticipated linear in 1n1-n behavior of $C_{v}$ from high $T$ expansion: energy fluctuations are dominated by uncorrelated spins and holes. Please refer to Appendix B to see how doping depends on chemical potential and temperature $T$.
@@
In this work, we set $h=0.1$ and $h=0.01$. The results of these fields match with each other quantitatively, proving that we are in the weak-field regime (see Appendix C). For weak-doping regime (e.g., $n\sim 0.95$), $\chi(T)$ develops a clear maximum at a temperature $T^{*}$ (see the inset of Fig. 2(B)), which tracks the build-up of strong but short-ranged AFM correlations: cooling below $T^{*}$ suppresses spin fluctuations as local singlets form; heating above $T^{*}$ melts the AFM background into a Curie-like regime. As doping increases, the maximum weakens and shifts to lower $T$. We thus identify $T^{*}$ as a pseudogap crossover scale set by short-range [34, 76]. Its evolution mirrors the suppression of the $C_{v}$ peak and anchors the microscopic reorganization discussed later. In Appendix D, we compare our iPEPS results for specific heat and magnetic susceptibility against METTS using matrix product states applied to cylinders and have appreciable agreement.
@@
Fig. 2(C) shows the isothermal compressibility $\kappa(\mu,T)=(\partial n/\partial\mu)_{T}$ together with filling $n(\mu)$ (inset). At high temperature, the response is nearly featureless, as expected for a classical mixture of spins and holes with weak correlations. Upon cooling to about $T\sim 0.3$ range, $\kappa$ develops a small peak as a function of $\mu$ (or equivalently $n$). This emerges in the same doping window where $\chi(T)$ exhibits a pronounced maximum. Two competing trends shape $\kappa$ here. Close to half filling, Mott constraints suppress charge fluctuations and reduce $\kappa$; moving away from half filling, the growth of short-range AFM domains lowers the energy cost of rearranging charge locally (e.g., by accommodating holes so as not to destroy the background AFM), mildly enhancing $\kappa$. The net result is a small but robust peak which grows as a function of decreasing temperature, consistent with a crossover — not a phase transition — in two dimensions at finite $T$. This intermediate-$T$ peak of $\kappa$ versus doping mirrors the cluster-DMFT Widom-line phenomenology of the doped Mott insulator, where $\kappa$ and related thermodynamic derivatives reach maxima above a pseudogap–metal critical endpoint and thus track the pseudogap boundary [38, 70, 69]. This phenomenon has also been tied to charge clustering [68] by one of the authors, and our METTS snapshots on a cylindrical lattice in Appendix E show the same growth of fluctuating hole clusters without macroscopic phase separation. There, we detail the analysis: we flag hole-rich sites via an adaptive threshold, group them into nearest-neighbor connected components on a cylinder of length $24$ and width $4$ with periodic boundary conditions, and compile density-weighted cluster-size histograms resolved by total hole mass. As the temperature is lowered, the probability for having clusters of a certain size shifts toward larger but still finite clusters. Our compressibility values also agree qualitatively with the finite-temperature Lanczos benchmarks of Ref. [30]. A recent Fermi–Hubbard quantum simulator study similarly mapped a pseudogap phase diagram via a doping-tuned compressibility maximum and additionally momentum-selective depletion of low-energy spectral weight, revealing a line of thermodynamic anomalies and suggesting a link to charge order [37].
IV Higher Order Correlators
@@
In Fig. 3, we evaluate the spin-spin-hole 3P correlators for bonds at distance $d=1$ or $\sqrt{2}$. In a Néel background the unconditional correlator $\langle B_{ab}\rangle$ is negative for nearest neighbours (d=1d{=}1) and positive for diagonals ($d{=}\sqrt{2}$). Hence a red (positive) $C^{(3)}$ on d=1d{=}1 means the AFM bond is weakened (made less negative) by the hole, whereas a blue (negative) $C^{(3)}$ on $d{=}\sqrt{2}$ means the originally positive diagonal correlation is suppressed, both indicating a local FM effect of the hole. At high to intermediate temperatures ($T\geq 1/4$), the correlation maps are short-ranged across $n=0.85$, $0.9$ and $0.95$, but grow upon cooling and with increasing filling, as shown in Fig. 3(A). At lower temperature $T=1/10$, the correlator extends over the whole $7\times 7$ window for all fillings with a much slower decay. Figs. 3(B), (C), and (D) show these general tendencies in more detail. Panels (B) and (C) quantify the spatial dependence of the 3P response at fixed $T/t=0.1$ as a function of the bond–hole separation $r$, for nearest-neighbor bonds ($d=1$) and diagonal bonds ($d=\sqrt{2}$), respectively. Upon increasing filling (reducing doping), the magnitude and the spatial range of the conditioned response increase, consistent with the growth of short-range AFM correlations in the background. Panel (D) shows the temperature evolution at fixed filling $n=0.85$: cooling extends the range of the 3P response and increases its magnitude, while the high-$T$ curves remain short-ranged. These panels make explicit that the dopant-conditioned spin response is primarily a short- to intermediate-range reorganization that strengthens markedly as AFM correlations build up on cooling.
@@
In Fig. 4 we contrast the single-dopant L-geometry $C^{(3)}_{L}$ with two different two-dopant conditionings: the well separated $T$ configuration $C^{(4)}_{T}$, and the nearest-neighbor plaquette configuration $C^{(4)}_{\parallel}$ as a function of the doping 1n1-n. For dopings $0.00\lesssim 1-n\lesssim 0.18$, both $C^{(3)}_{L}$ and $C^{(4)}_{T}$ are positive, then cross zero near $n=0.8$ and become negative upon further doping. The change in sign can be attributed to weakening of AFM correlations due to increased holes. Moreover, the near-additivity $C^{(4)}_{T}\simeq 2C^{(3)}_{L}$ holds on both sides of the crossing within the plotted scatter, indicating weak interaction between the well-separated dopants. Interpreting the sign relative to the background bond value, the positive regime signals a local ferromagnetic skewing around a dopant (AFM bonds made less negative), whereas beyond filling $n\sim 0.8$, the softened AFM background (reduced spin stiffness) leads the conditioned response to reinforce the residual AFM correlations (AFM bonds made more negative). By contrast, $C^{(4)}_{\parallel}$ is negative across the entire range and grows in magnitude away from half filling (increased doping 1n1-n) and on cooling, consistent with the following mechanism: two adjacent holes eliminate seven, rather than eight, JJ–links, lowering the exchange cost and cooperatively enforcing AFM on the parallel edge of the plaquette. This response, at finite $T$, provides a microscopic route toward incipient charge segregation marked by enhanced compressibility (as shown in Fig. 2(C)); upon further cooling, it can act as a precursor to stripe formation where holes line up to minimize exchange frustration while retaining kinetic energy. Finally, the doping where $C^{(3)}_{L}$ and $C^{(4)}_{T}$ change sign coincides with the regime where the uniform susceptibility peak fades in thermodynamics (as shown in Fig. 2(B)), reinforcing the link between the weakening spin stiffness and the reorganization of local spin textures.
@@
In Appendix F, we show results for two other kinds of correlators: the five-point correlator consisting of four spin operators on the nearest-neighbor sites surrounding a central hole projector, and the electron momentum distribution function.
V Discussion
In this work we provide thermodynamic-limit, finite-temperature benchmarks for the two-dimensional square-lattice ttJJ model down to $T/t\simeq 0.1$ and up to hole doping $1-n\lesssim 0.25$ using purification-iPEPS. Three main messages emerge. First, the bulk thermodynamics show a coherent crossover structure upon cooling: a broad spin-entropy feature in the specific heat $C_{v}(T)$ associated with the build-up of short-range antiferromagnetism, a pronounced maximum in the uniform susceptibility $\chi(T)$ at a scale $T^{\ast}$ that shifts and weakens with doping, and a shallow enhancement of the compressibility $\kappa(\mu,T)$ that develops upon cooling in a similar doping window. Second, dopant-conditioned correlators reveal the corresponding short-distance microscopic response: single holes behave as magnetic polarons that weaken nearby nearest-neighbor AFM bonds relative to the unconditional background, whereas nearest-neighbor hole pairs generate a cooperative response that reinforces antiferromagnetism on the plaquette edge parallel to the pair, consistent with a local reduction of exchange frustration when two holes move together. Third, over the same parameter window the $d$-wave pairing correlator remains strictly short-ranged, indicating that in the regime accessed here the dominant finite-$T$ reorganization is magnetic and local, preceding any growth of superconducting coherence. These findings are timely in light of recent quantum-gas-microscope studies of the doped Hubbard model, where a susceptibility saturation/maximum versus temperature and a compressibility maximum versus doping develop upon cooling and have been used to map out a pseudogap crossover from thermodynamics [7, 37]. Our results provide complementary thermodynamic-limit benchmarks for the ttJJ model and, crucially, supply a real-space mechanism for how such thermodynamic anomalies can arise from local spin-texture reorganization around dopants without requiring macroscopic phase separation at finite temperature. A key limitation is that our two-site checkerboard iPEPS ansatz enforces translation symmetry up to two sublattices and therefore cannot represent stripe order or other longer-ranged symmetry breaking. Accessing the stripe regime and its interplay with pairing will require larger unit cells. Within the experimentally relevant temperature and doping range explored here, the combined thermodynamic and conditional-correlation data support a simple conclusion: the finite-temperature dopant-induced reorganization of the ttJJ model is predominantly short-ranged and magnetic, while superconducting correlations remain short-ranged.
The data used in the main text of this manuscript can be found in Ref. [79]. The data used in the Appendix can be given upon request.
Acknowledgements.
Appendix A Numerical Details